If 730-nm and 660-nm light passes through two slits 0.63 mm apart, how far apart are the second-order fringes for these two wavelengths on a screen 1.2 m away? Express your answer to two significant figures and include the appropriate units.
Answer #1

The expression for fringe width is given by the expression

x=\frac{n \lambda D}{d}

n\rightarrow \text{order of the fringe }

\lambda\rightarrow\text{wavelength of the light}

D → Distance bewteen slit and screen

d\rightarrow\text{Distance bewteen the two slits}

The distance between the fringes for the 730-nm and 660-nm light can be found using

\Delta x=\frac{n \Delta\lambda\ D}{d}

=\frac{(2)(730-660)\times 10^{-9}\times 1.2}{0.63\times 10^{-3}}

=0.26\ \text{mm}